OpenStudy (anonymous):

Given that P = (5, 1) and Q = (13, 8), find the component form and magnitude of vector PQ.

OpenStudy (anonymous):

@iPwnBunnies

OpenStudy (ipwnbunnies):

'Vector PQ' This means it starts at Point P and ends at Point Q. If the vector is written in component form: <x,y> Then the x and y components will be the difference of the components of Points Q and P. For example: The 'x' component of Vector PQ will be Xq - Xp = 13 - 5 = 8

OpenStudy (anonymous):

so 8,7

OpenStudy (anonymous):

what about the magnitude?

OpenStudy (aum):

\(\large PQ = <8,7>, \text{ Magnitude of PQ = }\sqrt{8^2 + 7^2} = ?\)

OpenStudy (anonymous):

\[\sqrt{113}\]

OpenStudy (aum):

yes.

OpenStudy (anonymous):

thank you C;

OpenStudy (aum):

yw